Determine the molarity of each of the following solutions from its osmotic pressure at 25°C. Include the van't Hoff factor for the solution when the factor is given.

Π= 0.0286 atm for a solution of urea (CH4N2O):

Π= 0.698 atm for a solution of KI, i = 1.90:
1
Expert's answer
2018-09-14T05:43:11-0400
Answer:
П = MRTi
M = П / RTi
(0.0258 atm) / ((0.08206 L atm/K mol) x (25 + 273) K x 1) = 0.001055 mol/L

(0.606 atm) / ((0.08206 L atm/K mol) x (25 + 273) K x 1.90) = 0.01304 mol/L